• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name M. Rvachev

    Log entry time 08:11:53 on December 21,1999

    Entry number 31677

    keyword=Shift summary

    12:00 Waiting for beam (energy is being switched to nominal 4044 MeV)

    1:15 EQ3 quenched for no reason; brought it back and cycled to 1600A.

    3:15 Got beam

    3:25 Started ep. ep energy measurement was not successful because of ep DAQ
    problems (cannot write events to disk).

    5:40 Tuning beam back to achromatic mode

    6:15 MCC cannot deliver beam with absolute fast feedback in energy on. They can
    start delivering with absolute FF and then switch to relative, which they claim has
    the same effect as running all the time with absolute, provided that they "retune"
    beam energy after every trip. After contacting Arun we agreed to run in this mode for the time being.

    7:00Now MCC says that the tune with switching from absolute feedback to relative
    (and preserving absolute energy value) is impossible, since they get trips with absolute mode.

    7:10 Now it turns out that tonight's shift so far has been in ep beam tune.

    7:30 MCC has some problems now.

    7:40 Now we got achromatic tune, but no feedbacks. Taking data (1514).

    1496-1500 junk (cosmics with beam on and off)

    1501 - empty target run with stable 5uA current in ep mode - can be used to study background with beam on

    1502-1506 junk

    1507 12C pointing run (raster off) (ep tune)

    1508-1509 raster check

    Start He3 elastic normalization runs (beam in ep tune):

    1510 He3 elastic, raster 4*4 at MCC (we see it as dx=6mm, dy=4mm at supposedly
    badly calibrated spot++), beam 5uA, lost beam in the end of run
    1511 -''-, lost beam in the end, good run
    1512 junk
    1513 He3 elastic, raster 4*4, 5uA, good run, a few trips during run

    Start achromatic tune runs. All feedbacks are off.

    1514 He3 elastic, 5uA, good run
    1515 He3 elastic, 5uA, good run (trip in the end)
    1516 He3 elastic, 5uA